A race car accelerates uniformly from

18.5 m/s to 23.6m/s in 21seconds.

Determine the acceleration of the car

and the distance traveled.


1
Expert's answer
2022-05-24T13:56:39-0400
"a=\frac{\Delta v}{\Delta t}\\
=\frac{23.6-18.5}{21.0}=0.24\:\rm m/s^2"